# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 3. Batch Normalization
Batch Normalization normalizes the inputs of each layer to have zero mean and unit variance, which stabilizes and accelerates training. It also acts as a form of regularization by adding noise to the inputs during training [79].
1 class MLP_BatchNorm(nn.Module):
@@
4 self.flatten = nn.Flatten()
5 self.fc1 = nn.Linear(28*28, 128)
6 self.bn1 = nn.BatchNorm1d(128)
7 self.relu = nn.ReLU()
8 self.fc2 = nn.Linear(128, 64)
9 self.bn2 = nn.BatchNorm1d(64)
10 self.fc3 = nn.Linear(64, 10)
@@
1 class MLP_Regularized(nn.Module):
@@
4 self.flatten = nn.Flatten()
5 self.fc1 = nn.Linear(28*28, 256)
6 self.bn1 = nn.BatchNorm1d(256)
7 self.relu = nn.ReLU()
8 self.dropout = nn.Dropout(p=0.5)
9 self.fc2 = nn.Linear(256, 128)
10 self.bn2 = nn.BatchNorm1d(128)
11 self.fc3 = nn.Linear(128, 10)
@@
Convolutional Neural Networks (CNNs) are a specialized type of neural network designed for processing structured grid data, such as images. CNNs have revolutionized the field of image processing by efficiently capturing spatial and hierarchical features, making them ideal for tasks such as image classification, object detection, and segmentation [98]. In this chapter, we will explore the key components of CNNs, including convolution operations, pooling layers, and fully connected layers, which together enable CNNs to learn from visual data.
@@ 15.1 Convolution Operations and Convolutional Layers
At the heart of a CNN is the convolution operation, which is used to extract local features from an image. Instead of treating each pixel individually, the convolution operation allows the network to learn spatial patterns, such as edges, textures, and shapes, by applying filters (also called kernels) to small regions of the image [114].
@@ 15.1.1 Convolution Kernels and Feature Extraction
A convolution kernel (or filter) is a small matrix of weights that is applied to the image to detect specific features. The kernel slides across the input image, computing a weighted sum of the pixel values at each location. This process is known as a convolution and results in a new matrix called a feature map, which highlights the presence of the learned features (e.g., edges or corners) in the image [114].
@@
20conv = nn.Conv2d(in_channels=1, out_channels=1, kernel_size=3, padding=1, bias=False)
21conv.weight = nn.Parameter(kernel)
@@ 15.1.2 Multi-channel Convolution
In real-world applications, most images are in color and consist of multiple channels, such as red, green, and blue (RGB). Multi-channel convolution handles these images by applying separate filters to each channel and then combining the results [98].
@@
9conv_rgb = nn.Conv2d(in_channels=3, out_channels=1, kernel_size=3, padding=1, bias=False)
@@ 15.2 Pooling Layers
Pooling layers are used in CNNs to reduce the spatial dimensions of feature maps while retaining the most important information. Pooling helps to reduce the computational complexity of the network and makes the model more robust to small translations or distortions in the image [55].
@@ 15.2.1 Max Pooling and Average Pooling
8max_pool = nn.MaxPool2d(kernel_size=2)
@@
12avg_pool = nn.AvgPool2d(kernel_size=2)
@@ 15.3 Fully Connected Layers and Classification
After passing through several convolution and pooling layers, the feature maps are flattened and passed into a fully connected layer (also known as a dense layer). The fully connected layer serves as the decision-making part of the CNN, where it combines the learned features and assigns probabilities to different classes in classification tasks [12].
@@ 15.3.1 Fully Connected Layers
The fully connected layer takes the flattened feature maps as input and uses a set of weights and biases to compute the final output. This layer is typically followed by a Softmax function, which converts the raw scores into probabilities for each class [135].
@@
2class SimpleCNN(nn.Module):
@@
5 self.conv1 = nn.Conv2d(in_channels=1, out_channels=16, kernel_size=3, padding=1)
6 self.pool = nn.MaxPool2d(kernel_size=2)
7 self.fc1 = nn.Linear(16 * 14 * 14, 128) # Assuming input image size is 28x28
8 self.fc2 = nn.Linear(128, 10) # Output layer for 10 classes
@@ 15.5 Classic CNN Architectures
Over the years, Convolutional Neural Networks (CNNs) have evolved significantly, leading to the development of more powerful and complex architectures. These architectures have pushed the boundaries of computer vision tasks, from simple digit recognition to highly challenging tasks like object detection in large datasets. In this section, we will explore some of the most influential CNN architectures, including LeNet, AlexNet, VGGNet, and ResNet, providing detailed explanations of their structure, innovations, and applications [96].
@@ 15.5.1 LeNet
LeNet is one of the earliest CNN architectures, developed by Yann LeCun in the late 1980s and early 1990s. It was designed for handwritten digit recognition and became famous for its performance on the MNIST dataset, a dataset of handwritten digits from 0 to 9 [97].
@@
8 class LeNet(nn.Module):
@@
11 self.conv1 = nn.Conv2d(1, 6, kernel_size=5) # 1 input channel, 6 output channels, 5x5 kernel
12 self.pool = nn.AvgPool2d(2, 2) # 2x2 pooling
13 self.conv2 = nn.Conv2d(6, 16, kernel_size=5) # 16 filters
14 self.fc1 = nn.Linear(16 * 4 * 4, 120) # Fully connected layer
15 self.fc2 = nn.Linear(120, 84)
16 self.fc3 = nn.Linear(84, 10) # Output for 10 classes
@@ 15.5.2 AlexNet
AlexNet is a deeper and more complex CNN architecture that won the 2012 ImageNet Large Scale Visual Recognition Challenge (ILSVRC) [174]. Developed by Alex Krizhevsky, Ilya Sutskever, and Geoffrey Hinton, AlexNet introduced several innovations, such as the use of ReLU activations and dropout for regularization, to significantly improve performance [92].
@@
1 class AlexNet(nn.Module):
@@
4 self.conv1 = nn.Conv2d(3, 96, kernel_size=11, stride=4, padding=2)
5 self.pool = nn.MaxPool2d(3, 2)
6 self.conv2 = nn.Conv2d(96, 256, kernel_size=5, padding=2)
7 self.conv3 = nn.Conv2d(256, 384, kernel_size=3, padding=1)
8 self.conv4 = nn.Conv2d(384, 384, kernel_size=3, padding=1)
9 self.conv5 = nn.Conv2d(384, 256, kernel_size=3, padding=1)
10 self.fc1 = nn.Linear(256 * 6 * 6, 4096)
11 self.fc2 = nn.Linear(4096, 4096)
12 self.fc3 = nn.Linear(4096, 1000)
@@ 15.5.3 VGGNet
VGGNet was introduced by the Visual Geometry Group at the University of Oxford and achieved impressive results in the 2014 ImageNet challenge. VGGNet is known for its simplicity and uniformity, using small $3\times 3$ filters throughout the network [182].
@@
1 class VGGNet(nn.Module):
@@
4 self.conv_layers = nn.Sequential(
5 nn.Conv2d(3, 64, kernel_size=3, padding=1),
6 nn.ReLU(),
7 nn.Conv2d(64, 64, kernel_size=3, padding=1),
8 nn.ReLU(),
9 nn.MaxPool2d(2, 2), # Block 1
@@
11 nn.Conv2d(64, 128, kernel_size=3, padding=1),
12 nn.ReLU(),
13 nn.Conv2d(128, 128, kernel_size=3, padding=1),
14 nn.ReLU(),
15 nn.MaxPool2d(2, 2), # Block 2
@@
17 nn.Conv2d(128, 256, kernel_size=3, padding=1),
18 nn.ReLU(),
19 nn.Conv2d(256, 256, kernel_size=3, padding=1),
20 nn.ReLU(),
21 nn.MaxPool2d(2, 2) # Block 3
@@
23 self.fc_layers = nn.Sequential(
24 nn.Linear(256 * 8 * 8, 4096),
25 nn.ReLU(),
26 nn.Linear(4096, 4096),
27 nn.ReLU(),
28 nn.Linear(4096, 1000) # Output for 1000 classes
@@ 15.5.4 ResNet
ResNet, short for Residual Network, was introduced by Kaiming He and his colleagues in 2015. ResNet solved one of the biggest challenges in training deep networks—vanishing gradients—by introducing skip connections (or residual connections). This breakthrough enabled the training of networks with hundreds or even thousands of layers [67].
@@
A residual block is the core building block of ResNet. Instead of learning the direct mapping from the input $x$ to the output $y$, a residual block learns the residual, $F(x)=y-x$, which simplifies the learning process. The residual block has the following structure:
@@
1 class ResidualBlock(nn.Module):
@@
4 self.conv1 = nn.Conv2d(in_channels, out_channels, kernel_size=3, stride=stride, padding=1)
5 self.bn1 = nn.BatchNorm2d(out_channels)
6 self.conv2 = nn.Conv2d(out_channels, out_channels, kernel_size=3, stride=1, padding=1)
7 self.bn2 = nn.BatchNorm2d(out_channels)
8 self.shortcut = nn.Sequential()
@@
10 self.shortcut = nn.Sequential(
11 nn.Conv2d(in_channels, out_channels, kernel_size=1, stride=stride),
12 nn.BatchNorm2d(out_channels)
@@
22 class ResNet(nn.Module):
@@
25 self.conv1 = nn.Conv2d(3, 64, kernel_size=7, stride=2, padding=3)
26 self.bn1 = nn.BatchNorm2d(64)
27 self.pool = nn.MaxPool2d(3, 2)
@@
30 self.fc = nn.Linear(128 * 8 * 8, num_classes)
@@
37 return nn.Sequential(*layers)
@@ 16.1 CNN for Image Classification
8 class SimpleCNN(nn.Module):
@@
11 self.conv1 = nn.Conv2d(3, 16, 3, padding=1)
12 self.pool = nn.MaxPool2d(2, 2)
13 self.conv2 = nn.Conv2d(16, 32, 3, padding=1)
14 self.fc1 = nn.Linear(32 * 8 * 8, 128)
15 self.fc2 = nn.Linear(128, 10)
@@ 16.2 Transfer Learning
Transfer learning is a technique where a pre-trained model is fine-tuned for a new task. This approach is particularly useful when we have limited data for the new task, as pre-trained models have already learned valuable features from large datasets like ImageNet. Instead of training a CNN from scratch, transfer learning allows us to leverage the knowledge learned by models such as VGG, ResNet, or Inception [142].
@@
2 from torch import nn
@@
13 resnet.fc = nn.Linear(num_ftrs, 10)
@@ 16.3 Data Augmentation
Data augmentation is a technique used to artificially increase the size of the training dataset by creating modified versions of existing images. This helps improve the model’s generalization ability by exposing it to a wider variety of data during training, reducing overfitting [181].
@@
Image segmentation is a critical task in computer vision where the goal is to classify each pixel of an image into a specific category[161]. Unlike image classification, which assigns a label to the entire image, segmentation involves making pixel-level predictions, making it ideal for tasks such as medical imaging, autonomous driving, and scene understanding. Deep learning has significantly improved segmentation accuracy, especially with architectures like Fully Convolutional Networks (FCNs), U-Net, and the DeepLab series [161]. In this chapter, we explore these state-of-the-art methods in detail.
@@ 17.1 Fully Convolutional Networks (FCN)
A Fully Convolutional Network (FCN) is a type of neural network designed specifically for image segmentation tasks. Unlike traditional convolutional networks that use fully connected layers at the end for classification, FCNs replace fully connected layers with convolutional layers. This allows the network to produce spatial, pixel-level predictions instead of a single class label for the entire image. The key advantage of FCNs is that they can handle input images of arbitrary size and output a dense prediction map where each pixel is assigned a label [115].
@@
9 self.conv1 = nn.Conv2d(3, 64, kernel_size=3, padding=1)
10 self.conv2 = nn.Conv2d(64, 128, kernel_size=3, padding=1)
11 self.conv3 = nn.Conv2d(128, 256, kernel_size=3, padding=1)
@@
14 self.deconv1 = nn.ConvTranspose2d(256, 128, kernel_size=2, stride=2)
15 self.deconv2 = nn.ConvTranspose2d(128, 64, kernel_size=2, stride=2)
16 self.deconv3 = nn.ConvTranspose2d(64, num_classes, kernel_size=2, stride=2)

### v2

@@ 3. Batch Normalization
Batch Normalization normalizes the inputs of each layer to have zero mean and unit variance, which stabilizes and accelerates training. It also acts as a form of regularization by adding noise to the inputs during training [ioffe2015batch].
1 class MLP_BatchNorm($n$.Module):
@@
4 self.flatten = $n$.Flatten()
5 self.fc1 = $n$.Linear(28*28, 128)
6 self.bn1 = $n$.BatchNorm1d(128)
7 self.relu = $n$.ReLU()
8 self.fc2 = $n$.Linear(128, 64)
9 self.bn2 = $n$.BatchNorm1d(64)
10 self.fc3 = $n$.Linear(64, 10)
@@
1 class MLP_Regularized($n$.Module):
@@
4 self.flatten = $n$.Flatten()
5 self.fc1 = $n$.Linear(28*28, 256)
6 self.bn1 = $n$.BatchNorm1d(256)
7 self.relu = $n$.ReLU()
8 self.dropout = $n$.Dropout(p=0.5)
9 self.fc2 = $n$.Linear(256, 128)
10 self.bn2 = $n$.BatchNorm1d(128)
11 self.fc3 = $n$.Linear(128, 10)
@@
Convolutional Neural Networks (CNNs) are a specialized type of neural network designed for processing structured grid data, such as images. CNNs have revolutionized the field of image processing by efficiently capturing spatial and hierarchical features, making them ideal for tasks such as image classification, object detection, and segmentation [lecun2010convolutional]. In this chapter, we will explore the key components of CNNs, including convolution operations, pooling layers, and fully connected layers, which together enable CNNs to learn from visual data.
@@ 15.1 Convolution Operations and Convolutional Layers
At the heart of a CNN is the convolution operation, which is used to extract local features from an image. Instead of treating each pixel individually, the convolution operation allows the network to learn spatial patterns, such as edges, textures, and shapes, by applying filters (also called kernels) to small regions of the image [liu2018feature].
@@ 15.1.1 Convolution Kernels and Feature Extraction
A convolution kernel (or filter) is a small matrix of weights that is applied to the image to detect specific features. The kernel slides across the input image, computing a weighted sum of the pixel values at each location. This process is known as a convolution and results in a new matrix called a feature map, which highlights the presence of the learned features (e.g., edges or corners) in the image [liu2018feature].
@@
20conv = $n$.Conv2d(in_channels=1, out_channels=1, kernel_size=3, padding=1, bias=False)
21conv.weight = $n$.Parameter(kernel)
@@ 15.1.2 Multi-channel Convolution
In real-world applications, most images are in color and consist of multiple channels, such as red, green, and blue (RGB). Multi-channel convolution handles these images by applying separate filters to each channel and then combining the results [lecun2010convolutional].
@@
9conv_rgb = $n$.Conv2d(in_channels=3, out_channels=1, kernel_size=3, padding=1, bias=False)
@@ 15.2 Pooling Layers
Pooling layers are used in CNNs to reduce the spatial dimensions of feature maps while retaining the most important information. Pooling helps to reduce the computational complexity of the network and makes the model more robust to small translations or distortions in the image [gholamalinezhad2020pooling].
@@ 15.2.1 Max Pooling and Average Pooling
8max_pool = $n$.MaxPool2d(kernel_size=2)
@@
12avg_pool = $n$.AvgPool2d(kernel_size=2)
@@ 15.3 Fully Connected Layers and Classification
After passing through several convolution and pooling layers, the feature maps are flattened and passed into a fully connected layer (also known as a dense layer). The fully connected layer serves as the decision-making part of the CNN, where it combines the learned features and assigns probabilities to different classes in classification tasks [basha2020impact].
@@ 15.3.1 Fully Connected Layers
The fully connected layer takes the flattened feature maps as input and uses a set of weights and biases to compute the final output. This layer is typically followed by a Softmax function, which converts the raw scores into probabilities for each class [nwankpa2018activation].
@@
2class SimpleCNN($n$.Module):
@@
5 self.conv1 = $n$.Conv2d(in_channels=1, out_channels=16, kernel_size=3, padding=1)
6 self.pool = $n$.MaxPool2d(kernel_size=2)
7 self.fc1 = $n$.Linear(16 * 14 * 14, 128) # Assuming input image size is 28x28
8 self.fc2 = $n$.Linear(128, 10) # Output layer for 10 classes
@@ 15.5 Classic CNN Architectures
Over the years, Convolutional Neural Networks (CNNs) have evolved significantly, leading to the development of more powerful and complex architectures. These architectures have pushed the boundaries of computer vision tasks, from simple digit recognition to highly challenging tasks like object detection in large datasets. In this section, we will explore some of the most influential CNN architectures, including LeNet, AlexNet, VGGNet, and ResNet, providing detailed explanations of their structure, innovations, and applications [lecun2015deep].
@@ 15.5.1 LeNet
LeNet is one of the earliest CNN architectures, developed by Yann LeCun in the late 1980s and early 1990s. It was designed for handwritten digit recognition and became famous for its performance on the MNIST dataset, a dataset of handwritten digits from 0 to 9 [lecun1998gradient].
@@
8 class LeNet($n$.Module):
@@
11 self.conv1 = $n$.Conv2d(1, 6, kernel_size=5) # 1 input channel, 6 output channels, 5x5 kernel
12 self.pool = $n$.AvgPool2d(2, 2) # 2x2 pooling
13 self.conv2 = $n$.Conv2d(6, 16, kernel_size=5) # 16 filters
14 self.fc1 = $n$.Linear(16 * 4 * 4, 120) # Fully connected layer
15 self.fc2 = $n$.Linear(120, 84)
16 self.fc3 = $n$.Linear(84, 10) # Output for 10 classes
@@ 15.5.2 AlexNet
AlexNet is a deeper and more complex CNN architecture that won the 2012 ImageNet Large Scale Visual Recognition Challenge (ILSVRC) [russakovsky2015imagenet]. Developed by Alex Krizhevsky, Ilya Sutskever, and Geoffrey Hinton, AlexNet introduced several innovations, such as the use of ReLU activations and dropout for regularization, to significantly improve performance [krizhevsky2012imagenet].
@@
1 class AlexNet($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(3, 96, kernel_size=11, stride=4, padding=2)
5 self.pool = $n$.MaxPool2d(3, 2)
6 self.conv2 = $n$.Conv2d(96, 256, kernel_size=5, padding=2)
7 self.conv3 = $n$.Conv2d(256, 384, kernel_size=3, padding=1)
8 self.conv4 = $n$.Conv2d(384, 384, kernel_size=3, padding=1)
9 self.conv5 = $n$.Conv2d(384, 256, kernel_size=3, padding=1)
10 self.fc1 = $n$.Linear(256 * 6 * 6, 4096)
11 self.fc2 = $n$.Linear(4096, 4096)
12 self.fc3 = $n$.Linear(4096, 1000)
@@ 15.5.3 VGGNet
VGGNet was introduced by the Visual Geometry Group at the University of Oxford and achieved impressive results in the 2014 ImageNet challenge. VGGNet is known for its simplicity and uniformity, using small $3\times 3$ filters throughout the network [simonyan2014very].
@@
1 class VGGNet($n$.Module):
@@
4 self.conv_layers = $n$.Sequential(
5 $n$.Conv2d(3, 64, kernel_size=3, padding=1),
6 $n$.ReLU(),
7 $n$.Conv2d(64, 64, kernel_size=3, padding=1),
8 $n$.ReLU(),
9 $n$.MaxPool2d(2, 2), # Block 1
@@
11 $n$.Conv2d(64, 128, kernel_size=3, padding=1),
12 $n$.ReLU(),
13 $n$.Conv2d(128, 128, kernel_size=3, padding=1),
14 $n$.ReLU(),
15 $n$.MaxPool2d(2, 2), # Block 2
@@
17 $n$.Conv2d(128, 256, kernel_size=3, padding=1),
18 $n$.ReLU(),
19 $n$.Conv2d(256, 256, kernel_size=3, padding=1),
20 $n$.ReLU(),
21 $n$.MaxPool2d(2, 2) # Block 3
@@
23 self.fc_layers = $n$.Sequential(
24 $n$.Linear(256 * 8 * 8, 4096),
25 $n$.ReLU(),
26 $n$.Linear(4096, 4096),
27 $n$.ReLU(),
28 $n$.Linear(4096, 1000) # Output for 1000 classes
@@ 15.5.4 ResNet
ResNet, short for Residual Network, was introduced by Kaiming He and his colleagues in 2015. ResNet solved one of the biggest challenges in training deep networks—vanishing gradients—by introducing skip connections (or residual connections). This breakthrough enabled the training of networks with hundreds or even thousands of layers [he2016deep].
@@
A residual block is the core building block of ResNet. Instead of learning the direct mapping from the input $x$ to the output $y$, a residual block learns the residual, F(x)=yxF(x)=y-x, which simplifies the learning process. The residual block has the following structure:
@@
1 class ResidualBlock($n$.Module):
@@
4 self.conv1 = $n$.Conv2d(in_channels, out_channels, kernel_size=3, stride=stride, padding=1)
5 self.bn1 = $n$.BatchNorm2d(out_channels)
6 self.conv2 = $n$.Conv2d(out_channels, out_channels, kernel_size=3, stride=1, padding=1)
7 self.bn2 = $n$.BatchNorm2d(out_channels)
8 self.shortcut = $n$.Sequential()
@@
10 self.shortcut = $n$.Sequential(
11 $n$.Conv2d(in_channels, out_channels, kernel_size=1, stride=stride),
12 $n$.BatchNorm2d(out_channels)
@@
22 class ResNet($n$.Module):
@@
25 self.conv1 = $n$.Conv2d(3, 64, kernel_size=7, stride=2, padding=3)
26 self.bn1 = $n$.BatchNorm2d(64)
27 self.pool = $n$.MaxPool2d(3, 2)
@@
30 self.fc = $n$.Linear(128 * 8 * 8, num_classes)
@@
37 return $n$.Sequential(*layers)
@@ 16.1 CNN for Image Classification
8 class SimpleCNN($n$.Module):
@@
11 self.conv1 = $n$.Conv2d(3, 16, 3, padding=1)
12 self.pool = $n$.MaxPool2d(2, 2)
13 self.conv2 = $n$.Conv2d(16, 32, 3, padding=1)
14 self.fc1 = $n$.Linear(32 * 8 * 8, 128)
15 self.fc2 = $n$.Linear(128, 10)
@@ 16.2 Transfer Learning
Transfer learning is a technique where a pre-trained model is fine-tuned for a new task. This approach is particularly useful when we have limited data for the new task, as pre-trained models have already learned valuable features from large datasets like ImageNet. Instead of training a CNN from scratch, transfer learning allows us to leverage the knowledge learned by models such as VGG, ResNet, or Inception [pan2009survey].
@@
2 from torch import $n$
@@
13 resnet.fc = $n$.Linear(num_ftrs, 10)
@@ 16.3 Data Augmentation
Data augmentation is a technique used to artificially increase the size of the training dataset by creating modified versions of existing images. This helps improve the model’s generalization ability by exposing it to a wider variety of data during training, reducing overfitting [shorten2019survey].
@@
Image segmentation is a critical task in computer vision where the goal is to classify each pixel of an image into a specific category[ren2024deeplearningmachinelearning]. Unlike image classification, which assigns a label to the entire image, segmentation involves making pixel-level predictions, making it ideal for tasks such as medical imaging, autonomous driving, and scene understanding. Deep learning has significantly improved segmentation accuracy, especially with architectures like Fully Convolutional Networks (FCNs), U-Net, and the DeepLab series [ren2024deeplearningmachinelearning]. In this chapter, we explore these state-of-the-art methods in detail.
@@ 17.1 Fully Convolutional Networks (FCN)
A Fully Convolutional Network (FCN) is a type of neural network designed specifically for image segmentation tasks. Unlike traditional convolutional networks that use fully connected layers at the end for classification, FCNs replace fully connected layers with convolutional layers. This allows the network to produce spatial, pixel-level predictions instead of a single class label for the entire image. The key advantage of FCNs is that they can handle input images of arbitrary size and output a dense prediction map where each pixel is assigned a label [long2015fully].
@@
9 self.conv1 = $n$.Conv2d(3, 64, kernel_size=3, padding=1)
10 self.conv2 = $n$.Conv2d(64, 128, kernel_size=3, padding=1)
11 self.conv3 = $n$.Conv2d(128, 256, kernel_size=3, padding=1)
@@
14 self.deconv1 = $n$.ConvTranspose2d(256, 128, kernel_size=2, stride=2)
15 self.deconv2 = $n$.ConvTranspose2d(128, 64, kernel_size=2, stride=2)
16 self.deconv3 = $n$.ConvTranspose2d(64, num_classes, kernel_size=2, stride=2)